Determine the critical numbers of the given function and classify each critical point as a relative maximum, a relative minimum, or neither. f(t) =(t^2)/t^2+t-2

1
Expert's answer
2021-12-22T17:59:33-0500

"\displaystyle
f(t) = \frac{t^2}{t^2+t-2}\\
\text{Differentiating, we obtain}\\
f'(t)=\frac{t^2-4t}{(t^2+t-2)^2}\\
f'(t)=\frac{t^2-4t}{(t^2+t-2)^2}=0\\
\implies t = 0, t= 4\\
\text{Using the 2nd test derivative, we evaluate 2nd order derivative}\\
f''(t) =-\frac{2(t^3-6t^2-4)}{(t-1)^3(t+2)^3}\\
\text{When t = 0, $f''(t) =-1$, using the 2nd derivative test, f is maximum at t = 0}\\
\text{when t = 4, $f''(t) = 0.012$, using 2nd derivative test, f is minimum at t = 4}\\"